### v1

# PTR: A Pre-trained Language Model for Trajectory Recovery

Abstract.
Spatiotemporal trajectory data is vital for web-of-things services and is extensively collected and analyzed by web-based hardware and platforms. However, issues such as service interruptions and network instability often lead to sparsely recorded trajectories, resulting in a loss of detailed movement data. As a result, recovering these trajectories to restore missing information becomes essential. Despite progress, several challenges remain unresolved. First, the lack of large-scale dense trajectory data hampers the performance of existing deep learning methods, which rely heavily on abundant data for supervised training. Second, current methods struggle to generalize across sparse trajectories with varying sampling intervals, necessitating separate re-training for each interval and increasing computational costs. Third, external factors crucial for the recovery of missing points are not fully incorporated.
To address these challenges, we propose a framework called PTR. This framework mitigates the issue of limited dense trajectory data by leveraging the capabilities of pre-trained language models (PLMs). PTR incorporates an explicit trajectory prompt and is trained on datasets with multiple sampling intervals, enabling it to generalize effectively across different intervals in sparse trajectories. To capture external factors, we introduce an implicit trajectory prompt that models road conditions, providing richer information for recovering missing points. Additionally, we present a trajectory embedder that encodes trajectory points and transforms the embeddings of both observed and missing points into a format comprehensible to PLMs. Experimental results on two public trajectory datasets with three sampling intervals demonstrate the efficacy and scalability of PTR.
1. Introduction
A trajectory is represented as a sequence of timestamped locations that depict the movement of individuals and vehicles. With the advances in web-of-things services like ride-hailing and sharing bikes and web-based platforms like Google Maps and Mapbox in recent years, massive trajectory data is collected and analyzed. Such data plays a pivotal role in various applications, including urban planning (Wang et al., 2024), traffic management (Rehena and Janssen, 2018; Wang et al., 2021), and personalized location services (Zhao et al., 2020; Wang et al., 2020).
Collecting trajectory data involves the coordination of multiple devices (Chen et al., 2024). The process begins with continuously recorded GPS coordinates from devices like smartphones or IoT sensors at regular intervals. These data points are then sent to central servers or databases for storage and analysis. However, factors such as network transmission instability, device power-saving settings, and poor coverage in high-traffic or remote areas will result in the loss of trajectory points, making the collected data sparse. Such sparse trajectories failed to accurately describe movement behavior and route choices, thus limiting their effectiveness in trajectory-based applications. To mitigate this problem, it is essential to recover the missing points from sparse trajectories which improves the integrity of trajectory information. Such task is usually termed trajectory recovery.
Many works have been proposed to tackle trajectory recovery task, which can be divided into two categories: free space trajectory recovery (Chen et al., 2011, 2023a) and map-matched trajectory recovery (Wei et al., 2024; Chen et al., 2023b). Free space trajectory recovery aims to directly restore the coordinates of missing trajectory points but cannot ensure that the recovered trajectory is constrained to the road network. Typically, an additional map-matching process is required before these trajectories can be used in navigation-based applications. Map-matched trajectory recovery focuses on directly recovering trajectories on the road network. In this paper, we focus on the latter setting because it can be directly utilized by downstream tasks without the need for further processing. Despite the progress made by existing methods on map-matched trajectory recovery, they still face some challenges as listed below.
1) The availability of dense trajectories is often limited. Deep learning-based trajectory recovery models are inherently data-driven, requiring a substantial number of dense trajectories as labels for supervised training. However, due to the aforementioned factors, most collected and stored trajectories are sparse, resulting in a limited availability of large-scale dense trajectory data. As a result, the lack of sufficient dense trajectories for training makes existing deep learning models prone to overfitting and leads to diminished performance.
2) Existing models struggle to generalize to sparse trajectories with varying sampling intervals. Sparse trajectories often exhibit inconsistent sampling intervals due to network instability. However, most trajectory recovery models are trained on sparse trajectories with fixed sampling intervals, limiting their ability to generalize to trajectories with varying intervals. As a result, when confronted with trajectories that have different sampling intervals, these models require separate training for each interval to recover trajectories effectively, leading to significantly increased training costs.
Figure 1. An example of road network evolution process.
3) External factors related to the recovery of missing points are not incorporated. External factors, such as road conditions, provide valuable insights that can enhance trajectory recovery. For instance, in the scenario illustrated in Figure 1, given the observed points $p_{1}$ and $p_{4}$ and the road conditions at time $t_{1}$, it is more likely that the user took route $R_{1}$ rather than $R_{2}$. Additionally, based on the road conditions at times $t_{2}$ and $t_{3}$, as shown in Figures 1 (b) and (c), it becomes easier to infer that the user first accelerated and then decelerated between $p_{1}$ and $p_{4}$. However, existing methods only consider information from the known trajectory points, overlooking the significance of external factors.
To address the aforementioned challenges, we propose a Pre-trained Language Model for Trajectory Recovery (PTR). By leveraging a pre-trained language model (PLM) that has been trained on large-scale corpus datasets and fine-tuning it with a limited set of dense trajectories, PTR can effectively recover trajectories even when only scarce dense trajectory data is available. PTR incorporates two key components to generalize to sparse trajectories with varying sampling intervals: first, an explicit trajectory prompt, which uses natural language to describe both the trajectory recovery task and its characteristics. This activates the PLM’s ability to recover trajectories and recognize the trajectory’s sampling interval. Second, a joint training strategy, where PTR is trained on trajectory data with multiple sampling intervals, further enhancing its generalization capabilities. To account for external factors, we introduce an implicit trajectory prompt, which integrates rich contextual information, such as road conditions, to provide insights on missing points. Lastly, we propose a trajectory embedder that encodes both observed and missing trajectory points, transforming them into representations interpretable by PLMs. This allows the PLM to capture the spatiotemporal correlations within the trajectory.
Overall, the main contributions of our work can be summarized as follows:
  • We introduce PTR, a novel framework for trajectory recovery that utilizes PLMs to efficiently recover trajectories using a limited set of dense trajectories.
  • We propose explicit trajectory prompts that enable the PLM to accurately interpret the trajectory recovery task and determine the sampling intervals of the trajectories. Combined with a joint training strategy on trajectories with various sampling intervals, PTR generalizes effectively to sparse trajectories across different intervals.
  • We present an implicit trajectory prompt to model external factors influencing missing trajectory points and a trajectory embedder to transform the trajectory representation into a form that the PLM can process.
  • We conduct extensive experiments on two real-world trajectory datasets, each with three different sampling intervals. The results show that our proposed model achieves state-of-the-art performance in both accuracy and efficiency.
2. Related Work
2.1. Free-space Trajectory Recovery
Trajectory recovery in free space aims to directly restore the GPS coordinates of missing points. Early approaches relied on predefined rules to model mobile objects (Chen et al., 2016; Su et al., 2013). For example, Linear (Hoteit et al., 2014) assumes that vehicle movement follows uniform linear motion and employs linear interpolation for trajectory recovery. MPR (Chen et al., 2011) assumes vehicles travel along the most frequently used routes, utilizing historical trajectory data to determine route frequencies for recovery. Other studies (Banerjee et al., 2014; Zheng et al., 2012) use Markov models to capture spatial transitions between trajectory points. However, these methods are limited to modeling low-order transitions and struggle to capture the global spatial-temporal dependencies of trajectories.
Recent deep learning-based approaches have provided more effective solutions for trajectory recovery (Xu et al., 2017, 2022). DHTR (Wang et al., 2019) employs a GRU-based seq2seq model to analyze user route transitions, integrating Kalman filtering to improve recovery accuracy. AttnMove (Xia et al., 2021) uses attention mechanisms and LSTM to capture overall human movement preferences. PeriodicMove (Sun et al., 2021) models each user trajectory as a graph, employing graph neural networks to efficiently capture user preferences by leveraging historical route transitions. TERI (Chen et al., 2023a) introduces a two-stage framework based on transformers to recover trajectories with unequal spacing, while TrajBERT (Si et al., 2023) utilizes the BERT (Devlin et al., 2018) framework to recover implicit trajectories, focusing on coarser-grained outputs such as area or Wi-Fi IDs. Although these methods show promising results, they require an additional map-matching process to align the recovered trajectories with the road network before they can be applied in navigation systems. This two-stage process leads to error accumulation and introduces significant time overhead, as the map-matching step is time-consuming.
2.2. Map-matched Trajectory Recovery
Compared with the above methods, map-matched trajectory recovery incorporates the road network as input and aims to recover the trajectory directly onto the road. To this end, MTrajRec (Ren et al., 2021) first represents the trajectory point with a road segment and moving rate, utilizing the seq2seq-based multi-task framework to capture the spatiotemporal correlations within the trajectory. Following this idea, RNTrajRec (Chen et al., 2023b) further models the relationship between trajectories and road networks, designing a spatial-temporal transformer architecture to encode sparse trajectories. MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective, considering both micro and macro semantic information of trajectories.
Although these methods demonstrate promising progress in trajectory recovery, their performance are still hindered by the poor availability of large-scale dense trajectories and limited in generalization to different sampling intervals discussed in Section 1.
3. Preliminaries
3.1. Definitions
Definition 3.1 (Trajectory).
A trajectory is defined as a series of timestamped locations, denoted as $\mathcal{T}=\langle p_{1},\cdots,p_{|\mathcal{T}|}\rangle$ where $p_{i}=(lat_{i},lng_{i},t_{i})$ represents the latitude and longitude coordinates of an object at the time $t_{i},i\in\{1,\cdots,|\mathcal{T}|\}$. $|\mathcal{T}|$ is the length of trajectory. The sampling interval of the trajectory $\mathcal{T}$ is $t_{i}-t_{i-1},i\in\{2,\cdots,|\mathcal{T}|\}$.
Definition 3.2 (Road Network).
A road network is defined as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes. Each node $v\in\mathcal{V}$ signifies an intersection that links various road segments, and each node possesses attributes of latitude and longitude. $\mathcal{E}$ represents the set of edges, where each edge $e\in\mathcal{E}$ denotes road segment that connecting two intersections. An edge can be characterized by its starting intersection $e.start\in\mathcal{V}$ and ending intersection $e.end\in\mathcal{V}$.
Definition 3.3 (Map-matched Trajectory).
By employing the map-matching algorithm, a trajectory $\mathcal{T}$ can be projected onto the road network and return the map-matched trajectory $\mathcal{T}_{m}$. As a result, each point of the map-matched trajectory aligns accurately with a particular road. The map-matched trajectory is represented as $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$, where each $q_{j}=(e_{j},r_{j},t_{j})$ signifies the vehicle’s location at time $t_{j},j\in\{1,\cdots,|\mathcal{T}_{m}|\}$. Here, $e\in\mathcal{V}$ is the matched road segment, and $r$ is the moving ratio, which quantifies the proportion of the distance traveled from the starting point of a road segment in relation to the total length of that road segment.
Figure 2. An illustration of map-matched trajectory, road segment $e$, and moving ratio $r$.
Example 3.4.
Figure 2 gives an example of a map-matched trajectory. For a sparse trajectory $\mathcal{T}_{s}=\langle p_{1},p_{4},p_{7}\rangle$ with the sampling interval of 90 seconds, its corresponding dense map-matched trajectory is represented by $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{7}\rangle$ with a sampling interval of 30 seconds. Take a matching point $q_{6}$ as an example. It falls on the road $e_{3}$, so its road segment is $e_{3}$. The distance of $q_{6}$ from the start point of the road is represented as $d_{6}$, while the total length of $e_{3}$ is denoted by $l_{3}$. Therefore, the moving ratio $r_{6}$ can be calculated as $r_{6}=\frac{d_{6}}{l_{3}}$.
Figure 3. The framework of PTR, consists of Trajectory Prompt, Trajectory Embedder, and PLM-based Trajectory Encoder.
3.2. Problem Statement
Map-matched Trajectory Recovery. Given a sparse trajectory $\mathcal{T}_{s}=\langle p_{1},\cdots,p_{|\mathcal{T}_{s}|}\rangle$ with a sampling interval of $\mu$. The goal of map-matched trajectory recovery is to reconstruct the dense map-matched trajectory $\mathcal{T}_{m}=\langle q_{1},\cdots,q_{|\mathcal{T}_{m}|}\rangle$ with a sampling interval of $\epsilon$. Note that the sampling interval $\mu$ is larger than $\epsilon$.
4. Methodology
4.1. Overall Pipeline
In this paper, we present PTR, which fine-tunes a PLM pre-trained on a large-scale corpus to efficiently recover trajectories with limited dense data. The framework of PTR, shown in Figure 3, comprises three main components: trajectory prompts, a trajectory embedder, and a PLM-based trajectory encoder.
Trajectory prompts offer essential information from two key components. First, explicit trajectory prompts, expressed in natural language, outline the trajectory recovery task and key trajectory characteristics. These prompts activate the PLM’s ability to recover trajectories and identify varying sampling intervals in sparse trajectories. With joint training on trajectories across different intervals, PTR achieves robust generalization. Second, implicit trajectory prompts model road conditions, providing crucial external factors of missing points.
The trajectory embedder effectively enhances the PLM’s understanding of trajectories. It begins by normalizing sparse trajectories with varying sampling intervals $\mu$ into a uniform interval $\epsilon$. Next, each trajectory point, including both observed and missing points, is encoded into an embedding that captures both local and global semantics. Finally, the embedding is transformed into a format understood with the PLM.
In the PLM-based trajectory encoder, most of the encoder’s parameters are frozen to preserve the pre-trained knowledge, while the multi-head attention layer remains trainable, allowing the model to capture the spatiotemporal correlations within the trajectory. Finally, the output layer predicts the road segment $e$ and the moving ratio $r$ for the recovered map-matched trajectory at each step.
4.2. Trajectory Prompts
To enable PLMs to recover trajectories with different intervals and efficiently model external factors for missing trajectory points, we introduce explicit and implicit trajectory prompts.
4.2.1. Explicit Trajectory Prompt
The explicit trajectory prompt offers a structured textual description at both the task and trajectory levels, enabling PTR to adapt to varying sampling intervals. At the task level, it includes the following components:
  • ¡Task Part¿: Sparse trajectory recovery.
  • ¡Target Part¿: Output the road segment and moving ratio for each point in the trajectory.
[…]
At the trajectory level, the explicit trajectory prompt provides detailed characteristics of the trajectory, consisting of:
[…]
The ¡Time Part¿ provides the trajectory’s specific start and end times, helping the PLM understand the duration and potential time patterns, such as morning or evening peaks. The ¡Movement Part¿ supports the PLM in inferring the trajectory’s movement.
After obtaining the prompt for each part, we use the tokenizer and word token embedding in PLMs to convert text into embeddings. The embedding of ¡Task Part¿, ¡Target Part¿, ¡Content Part¿, ¡Time Part¿, and ¡Movement Part¿ are denoted by $\textbf{H}_{ta}$, $\textbf{H}_{tar}$, $\textbf{H}_{c}$, $\textbf{H}_{t}$, and $\textbf{H}_{m}$. These embeddings are then concatenated to form the overall explicit trajectory prompt embedding.
(1) $\textbf{H}^{e}=\textbf{H}_{ta}||\textbf{H}_{tar}||\textbf{H}_{c}||\textbf{H}_{t}||\textbf{H}_{m},$
where ||||| | is the concatenate operation.
4.2.2. Implicit Trajectory Prompt
We incorporate road conditions as external factors for missing trajectory points, as they provide insights into both the surrounding environment and the object’s movement, which are all helpful for trajectory recovery. For instance, vehicles typically slow down in congested areas and accelerate in less congested regions. Given the complexity and variability of real-world road conditions, which are difficult to express in natural language, we represent them as an implicit trajectory prompt.
To obtain the road conditions for each trajectory point, we first calculate the average road conditions across all areas and time intervals, then extract the relevant information for each point. Specifically, we divide the area of interest into a spatial grid of $I\times J$ cells and partition the time into $T$ slices. For each grid cell, we compute the total traffic, resulting in a regional flow matrix $\textbf{RC}\in\mathbb{R}^{I\times J\times T}$. To capture the spatiotemporal correlations among different regions, we apply a 2D convolution (2D CNN) along the spatial dimensions and a 1D CNN along the temporal dimension, yielding the road condition representation $\textbf{H}_{rc}\in\mathbb{R}^{I\times J\times T\times F}$, where $F$ is the number of features. Formally:
(2) $\mathbf{H}_{rc}=\text{Conv1d(Conv2d({RC}))}\in\mathbb{R}^{I\times J\times T\times F}$
Obviously, the road conditions of a trajectory point are indexed by its longitude, latitude, and timestamp. However, it is challenging for missing points as their geographic coordinates are unknown. To address this, we introduce a road condition diffusion strategy that approximates the missing conditions using neighboring traffic. The implementation details will be elaborated in Section 4.3.2.
Figure 4. An illustration of trajectory preprocessing layer.
4.3. Trajectory Embedder
To efficiently encode trajectory points and enable PLMs to better understand them, we introduce the trajectory embedder.
4.3.1. Trajectory Preprocessing Layer
For different sparse trajectories, their sampling intervals $\mu$ are not constant, such as 4 minutes, 2 minutes, or 1 minute. To convert these into a uniform sampling interval, we introduce a placeholder ‘[m]’ to mark missing trajectory points and generate a preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, where the length of $|\mathcal{T}^{\epsilon}|=\frac{p_{|\mathcal{T}_{s}|}.t-p_{1}.t}{\epsilon}$. Although the location of ‘[m]’ is unknown, its timestamp can still be calculated, as illustrated in Figure 4.
4.3.2. Trajectory Feature Extractor
To embed trajectory points, we propose the trajectory feature extractor. Given the preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, there are two cases for its trajectory point characteristics:
  • Case 1: the trajectory point $s\in\mathcal{T}^{\epsilon}$ is observed. i.e., $\exists\,k\in\{1,\cdots,|\mathcal{T}_{s}|\},p_{k}.t=s.t$.
  • Case 2: the location of trajectory point $s\in\mathcal{T}^{\epsilon}$ is missing. i.e., $s=[m]$, note that its time $s.t$ is known.
For case 1, we use Learnable Fourier Features (Tancik et al., 2020; Li et al., 2021) to encode the continuous latitude and longitude of $s$ into an $F$-dimensional vector using the feature mapping function $\Phi(x):\mathbb{R}\to\mathbb{R}^{F}$. Formally:
(3) $\Phi(x)=W_{\Phi}[\mathrm{cos}\text{ }xW_{r}||\mathrm{sin}\text{ }xW_{r}],$
where $W_{r}\in\mathbb{R}^{F/2}$ and $W_{\Phi}\in\mathbb{R}^{F\times F}$ represent the learnable parameters, and $x\in\{s.lat,s.lon\}$.
Using the feature mapping function $\Phi(\cdot)$, the relative information between points $x$ and $y$ can be captured through multiplication operations. This relative information helps infer crucial details like the distance between trajectory points, which is important for understanding movement. For instance, a larger distance between two points may indicate a higher likelihood of vehicle acceleration.
Additionally, considering that vehicles move within the road network, the relationship between a trajectory point and its surrounding road segments is critical. To measure the relationship between a trajectory point $s$ and a road segment $l$, we define a function $f(d_{s,l})$ based on their shortest distance:
(4) $f(d_{s,l})=\begin{cases}e^{-(\frac{d_{s,l}}{\kappa})^{2}}&\text{ if }d_{s,l}<\varphi_{dist},\\ 0&\text{ otherwise ,}\end{cases}$
where $d_{s,l}$ is the shortest distance between $s$ and road $l$, $\kappa$ is a hyperparameter, and $\varphi_{dist}$ is a distance threshold. Then, we obtain the road network representation $\mathbf{h}_{s}^{\text{road}}$ by:
(5) $\mathbf{h}_{s}^{\text{road}}=\frac{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})*\textbf{M}_{l}}{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})},$
where $\textbf{M}_{l}\in\mathbb{R}^{F}$ is the learnable embedding of road segment $l$, and $\mathcal{|V|}$ is the total number of road segments. Finally, the complete representation of a trajectory point $s$ is obtained as:
(6) $\mathbf{h}_{s}=W_{1}[(\Phi(s.lat)+\Phi(s.lng))||\mathbf{h}_{s}^{\text{road}}]+b_{1},$
where $W_{1}\in\mathbb{R}^{2F\times F}$ and $b_{1}\in\mathbb{R}^{F}$ is the learnable parameters.
For case 2, where the location of point $s$ is unknown, here we use road conditions to represent its features, as described in Section 4.2.2. Since the road conditions at a specific point are influenced by the surrounding conditions, which propagate along both temporal and spatial dimensions, we propose a road condition diffusion strategy to model the road conditions of the missing location.
First, for the missing point $s$, we identify the observed forward and backward trajectory points, $s_{f}$ and $s_{b}$, respectively. For instance, in Figure 4, suppose we have a missing point $s$ timestamped at 8:00:45, its observed forward point $s_{f}=p_{1}$ and the backward point $s_{b}=p_{2}$. Then, we retrieve the road conditions of $s_{f}$ and $s_{b}$ as follows:
(7) $\begin{split}\mathbf{h}_{rc}^{f}&=\mathbf{H}_{rc}[\pi_{lat}(s_{f}.lat),\pi_{lng}(s_{f}.lng),\pi_{t}(s_{f}.t)],\\ \mathbf{h}_{rc}^{b}&=\mathbf{H}_{rc}[\pi_{lat}(s_{b}.lat),\pi_{lng}(s_{b}.lng),\pi_{t}(s_{b}.t)],\end{split}$
where $\pi_{lat}$, $\pi_{lng}$, and $\pi_{t}$ are index functions mapping latitude, longitude, and time to their respective indices. Next, we calculate the time interval between $s$ and $s_{f}$ and $s_{b}$, denoted as $\triangle t_{f}=s.t-s_{f}.t$ and $\triangle t_{b}=s_{b}.t-s.t$. The road condition of point $s$ is then computed as:
(8) $\mathbf{h}_{rc}^{s}=\frac{e^{-\triangle t_{f}}\mathbf{h}_{rc}^{f}+e^{-\triangle t_{b}}\mathbf{h}_{rc}^{b}}{e^{-\triangle t_{f}}+e^{-\triangle t_{b}}}$
In addition, we encode the time intervals $\triangle t_{f}$ and $\triangle t_{b}$ to quantify the relative position of $s$. The final representation of $s$ is given by:
(9) $\begin{split}\mathbf{h}_{s}&=W_{2}[\textbf{m}\text{ }||\text{ FC}(\triangle t_{f}||\triangle t_{b})\text{ }||\text{ }\mathbf{h}_{rc}^{s}]+b_{2},\\ \text{FC}(\cdot)&:\mathbb{R}^{2}\to\mathbb{R}^{F},\end{split}$
where $\textbf{m}\in\mathbb{R}^{F}$ is a learnable vector to represent the location is missing, and FC()($\cdot$)( ⋅ ) is the fully connection layer. $W_{2}\in\mathbb{R}^{3F\times F}$ and $b_{2}\in\mathbb{R}^{F}$ are the learnable parameters. Finally, the overall trajectory representation is $\textbf{H}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times F}$.
4.3.3. Trajectory Feature Transformation
To enhance the model’s ability to comprehend trajectory features, we introduce $K$ reference tokens $\textbf{E}_{w}\in\mathbb{R}^{K\times F}$, inspired by (Lee et al., 2019; Guo et al., 2023), to bridge the connection between PLMs and the trajectory. These reference tokens are designed to capture the global semantics of the trajectory.
Given the trajectory embedding H, we first apply a 1D CNN to aggregate neighboring information and capture local movement patterns:
(10) H=Conv1d($\textbf{H})$ = Conv1d ( H )
Next, we compute self-attention between the trajectory embedding $\textbf{H}^{\prime}$ and the reference tokens $\textbf{E}_{w}$ to capture the global semantics, where $\textbf{H}^{\prime}$ acts as the query, and $\textbf{E}_{w}$ as both the key and value:
(11) $\textbf{H}^{\prime\prime}=\text{Attention}(\textbf{H}^{\prime},\textbf{E}_{w},\textbf{E}_{w})$
After transforming the embedding $\textbf{H}^{\prime\prime}$ in the token space, we concatenate it with the explicit trajectory prompt embedding $\textbf{H}^{e}$ to form the final trajectory representation:
(12) $\textbf{Z}=\textbf{H}^{e}||\textbf{H}^{\prime\prime}$
Then, we add the Transformer positional embedding $PE$ in each element of Z and feed it into PLMs encoder to encode trajectory.
4.4. PLM-based Trajectory Encoder
As shown in Figure 3, we use the stacked Transformer layers as the fundamental architecture of PLMs. To make PLMs fully adaptable to the trajectory recovery task, we employ the Low-Rank Adaptation (LoRA) algorithm (Hu et al., 2021). LoRA introduces additional parameters within the Transformer block, reducing the number of trainable parameters and computational complexity while enhancing the PLM’s performance on new tasks.
Specifically, in each self-attention block of the Transformer, we introduce additional weights $\Delta\textbf{W}_{q}\in\mathbb{R}^{F\times F}$, $\Delta\textbf{W}_{k}\in\mathbb{R}^{F\times F}$, and $\Delta\textbf{W}_{v}\in\mathbb{R}^{F\times F}$ for the query, key, and value matrices $\textbf{W}_{q}$, $\textbf{W}_{k}$, and $\textbf{W}_{v}$. To reduce the parameter count, the additional weights $\Delta\textbf{W}$ are decomposed into two low-rank matrices, $\textbf{B}\in\mathbb{R}^{F\times r}$ and $\textbf{C}\in\mathbb{R}^{r\times F}$, where $r\ll F$. The modified query, key, and value matrices in each self-attention block are updated as $\textbf{W}_{q}=\textbf{W}_{q}+\Delta\textbf{W}_{q}$, $\textbf{W}_{k}=\textbf{W}_{k}+\Delta\textbf{W}_{k}$, and $\textbf{W}_{v}=\textbf{W}_{v}+\Delta\textbf{W}_{v}$. Formally, the output of the $l$-th transformer layer is defined as:
(13) $\textbf{Z}^{l}=\text{LoRA(}\text{Transformer }(\textbf{Z}^{l-1}))$
After stacking $L$ Transformer layers, we discard the explicit trajectory prompt portion and obtain the output embeddings $\mathbf{Z}^{\prime}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times d}$.
For the $j$-th element of $\textbf{Z}^{\prime}$, we apply the softmax function to calculate the probability of each road segment and use the argmax function to return the predicted road segment $e_{j}$. We employ a MLP with a Sigmoid activation function to determine the moving ratio $r_{j}$.
4.5. Training
4.5.1. Loss Function
We use multi-task learning to simultaneously optimize road segment recovery and moving ratio recovery tasks. For the road segment recovery, we use the cross-entropy as the loss function:
(14) $\mathcal{L}_{e}=-\frac{1}{|\mathcal{T}^{\epsilon}|}\sum_{j=1}^{|\mathcal{T}^{\epsilon}|}\mathrm{log}(\hat{e}_{j}|\textbf{Z}^{\prime}_{j})$
For the moving ratio, we use the mean squared error as the loss function:
(15) $\mathcal{L}_{r}=\frac{1}{|\mathcal{T}^{\epsilon}|}\sum_{j=1}^{|\mathcal{T}^{\epsilon}|}|r_{j}-\hat{r}_{j}|^{2}$
The final loss function is $\mathcal{L}=\mathcal{L}_{e}+\lambda\mathcal{L}_{r}$, where $\lambda$ is a hyper-parameter to balance two tasks.
4.5.2. Training Strategy
To generalize PTR in recovering sparse trajectories with various sampling intervals, we propose the following training approach.
For the dense trajectory dataset $\mathbb{T}$, which is set as a 15-second sampling interval in our experiments. To create sparse trajectory datasets, we re-sample $\mathbb{T}$ at three different intervals: 1 minute, 2 minutes, and 4 minutes, resulting in three datasets: $\mathbb{T}_{1}$, $\mathbb{T}_{2}$, and $\mathbb{T}_{4}$. These are then combined to form a larger training dataset, $\mathbb{T}_{all}$.
(16) $\mathbb{T}_{all}=\mathbb{T}_{1}\cup\mathbb{T}_{2}\cup\mathbb{T}_{4}.$
By training PTR on $\mathbb{T}_{all}$, we achieve balanced performance across the different sampling intervals. Subsequently, we fine-tune the model on each individual sparse trajectory dataset to further improve trajectory recovery performance.
5. Experiments
In this section, we present comprehensive experiments to evaluate the effectiveness of PTR on two real-world trajectory datasets with three different sparse trajectory sampling intervals. These experiments are designed to answer the following key questions:
  • RQ1: How does PTR’s performance compare to other baseline methods across different sampling intervals of sparse trajectories?
  • RQ2: How well does PTR perform when trained with limited data?
  • RQ3: What is the contribution of each component in PTR to the overall model performance?
5.1. Datasets
We assess our model using two publicly available trajectory datasets from Chengdu, China, and Porto, Portugal. The Chengdu dataset [^1] https://outreach.didichuxing.com/ contains trajectory data collected in November 2016. The Porto dataset [^2] https://www.kaggle.com/competitions/pkdd-15-predict-taxi-service-trajectory-i/data includes trajectory from 442 taxis, collected from July 2013 to June 2014. Both datasets are standardized to a sampling interval of 15 seconds. We remove trajectories with travel times less than 5 minutes or exceeding 1 hour, as well as outlier trajectories. The road network data is obtained from the OpenStreetMap website [^3] http://www.openstreetmap.org/. We employ a map-matching algorithm (Newson and Krumm, 2009) to project the trajectory onto the road network and obtain the ground truth of the road segments and moving rates. The detailed overview of the dataset characteristics is provided in Table 4.
5.2. Baselines
To evaluate the effectiveness of our model, we compare PTR with ten baseline methods. These include five free space trajectory recovery models: HMM (Newson and Krumm, 2009) + ShortestPath, Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009), MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009), DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009), and AttnMove (Xia et al., 2021) + Rule, as well as five map-matched trajectory recovery models: MTrajRec (Ren et al., 2021), T2vec (Li et al., 2018) + Decoder, T3s (Yang et al., 2021) + Decoder, RNTrajRec (Chen et al., 2023b), and MM-STGED (Wei et al., 2024). The details of each baseline are provided in Appendix B.
5.3. Evaluation Metrics
We employ five common metrics to assess the effectiveness of our model followed (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024). For road segment recovery, we utilize Accuracy (Acc), Recall, and Precision (Prec) metrics to evaluate. A higher result indicates a more precise road segment recovery. And employ Mean Absolute Error (MAE) and the Root Mean Square Error (RMSE) in assessing the recovered GPS coordinates. Lower values of these metrics indicate a higher level of accuracy in the recovery results. The details of each metric are provided in Appendix C.
Table 1. Performance comparison on two datasets with sampling intervals at 4 minutes, 2 minutes, and 1 minute, respectively.
The best results are highlighted in bold, while the underline indicates the second-best results.
Sampling Interval Methods Chengdu Porto
[…]
$\mu=4$ minutes $\to\epsilon=15$ seconds HMM + ShortestPath 26.85 28.64 29.55 939.3 1047.7 20.19 26.22 33.51 886.9 941.5
Linear + HMM 26.42 30.45 36.15 974.5 1145.4 32.23 36.09 49.80 489.3 637.3
MPR + HMM 36.93 38.62 44.53 821.9 914.1 32.22 38.67 48.07 534.0 700.2
DHTR + HMM 41.48 57.34 50.48 673.6 911.3 32.02 58.29 45.61 456.7 627.1
AttnMove + Rule 63.43 73.97 78.72 358.2 916.7 49.31 48.62 78.03 310.0 621.3
MTrajRec 65.79 75.14 78.42 315.1 904.4 52.36 60.39 77.28 266.1 590.1
T3s + Decoder 65.60 75.26 78.14 318.2 926.3 52.24 60.24 77.80 270.4 594.9
T2vec + Decoder 66.51 75.68 78.27 307.5 915.2 53.13 60.27 77.62 256.3 571.0
RNTrajRec 67.66 75.59 79.97 306.1 886.0 54.59 60.42 79.20 248.1 549.1
MM-STGED 70.64 76.04 81.63 266.2 829.7 57.30 59.48 80.21 222.8 510.4
PTR 74.12 79.63 86.46 262.8 483.0 57.61 59.15 82.19 200.9 376.9
PTR + FT 74.58 80.09 86.63 253.2 465.7 57.67 59.08 82.05 200.8 370.2
$\mu=2$ minutes $\to\epsilon=15$ seconds HMM + ShortestPath 33.85 47.89 48.31 754.1 826.2 27.30 40.05 46.00 647.3 747.0
Linear + HMM 43.78 45.35 48.77 816.9 1054.7 49.35 50.45 63.87 408.7 609.8
MPR + HMM 49.88 54.62 50.94 474.8 899.0 49.76 52.97 62.86 409.8 610.9
DHTR + HMM 47.17 60.16 51.73 662.0 912.2 43.76 65.25 52.10 385.3 578.0
AttnMove + Rule 71.98 77.42 80.67 291.1 764.8 61.39 60.90 82.98 213.4 468.5
MTrajRec 74.52 78.25 81.09 254.5 885.7 61.65 65.65 78.99 179.9 451.5
T3s + Decoder 74.62 78.95 81.79 242.2 857.5 61.75 65.53 79.14 181.3 461.2
T2vec + Decoder 75.69 78.86 81.68 231.6 783.6 62.24 65.77 78.97 173.9 438.0
RNTrajRec 75.80 79.35 81.86 218.5 757.0 63.39 65.84 79.25 171.3 433.9
MM-STGED 78.14 80.06 83.58 197.2 696.0 65.69 66.15 80.74 152.5 400.8
PTR 81.76 84.31 88.38 187.4 366.2 66.40 66.52 82.14 141.9 294.6
PTR + FT 82.29 84.59 88.72 181.2 349.0 66.72 66.87 82.38 141.7 293.5
$\mu=1$ minute $\to\epsilon=15$ seconds HMM + ShortestPath 35.92 67.92 60.16 529.7 638.2 34.75 48.46 48.67 527.2 659.3
Linear + HMM 68.59 65.66 66.67 707.4 1005.2 66.17 64.72 75.22 368.3 571.1
MPR + HMM 62.25 62.67 60.53 418.8 659.0 66.27 65.66 74.51 402.6 628.2
DHTR + HMM 51.09 63.40 50.14 584.7 750.4 52.98 69.60 57.17 420.4 625.8
AttnMove + Rule 79.60 81.55 82.75 194.4 752.6 72.07 69.59 80.41 156.6 360.7
MTrajRec 81.12 81.73 83.75 187.1 718.4 71.65 70.92 80.35 114.9 332.3
T3s + Decoder 80.90 82.78 83.15 187.1 713.0 71.78 71.61 80.16 110.1 328.0
T2vec + Decoder 81.69 81.90 83.88 185.6 714.1 71.86 71.10 80.48 114.2 334.7
RNTrajRec 81.88 82.09 84.84 177.9 702.5 72.31 71.88 80.57 110.3 325.7
MM-STGED 84.26 84.15 85.92 154.0 633.5 73.16 72.27 80.81 108.2 321.9
PTR 87.17 87.99 90.52 141.4 290.8 74.42 72.78 82.82 95.6 211.7
PTR + FT 88.15 88.82 91.04 138.6 289.1 75.35 73.87 83.28 95.2 220.3
5.4. Settings
We split the trajectory dataset into training, validation, and testing sets in a 7:2:1 ratio. Following previous work (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024), for dense map-matched trajectories, we set the sampling interval $\epsilon=15$ seconds. To generate sparse trajectories, for each dense trajectory, we retain the first and last points and create three sparse versions by setting the sampling intervals $\mu=4$ minutes, 2 minutes, and 1 minute, respectively.
During training, we integrate sparse trajectories with three different sampling intervals to train PTR. After training, we evaluate its performance in recovering trajectories across various sampling intervals to assess its generalizability. To further enhance performance, we fine-tune the trained model for each specific sampling interval, allowing for a more precise recovery.
We employ PyTorch (Paszke et al., 2019) framework to implement PTR, with a learning rate of 1e-4 and a batch size of 64. BERT (Devlin et al., 2018) is selected as the foundation model for PLMs, as its encoder-only structure is well-suited to the reconstruction task due to its effective use of bi-directional contextual information from the trajectory (Ren et al., 2024). For road condition extraction, the area of interest is divided into a $64\times 64$ grid, and the time dimension is partitioned into hourly intervals. The hidden state dimension is set to $F=512$. In Equation 4, we set $\kappa=15$ and $\varphi_{dist}=50$ meters. The model is trained for 50 epochs with early stopping, using a patience of 10 epochs. All experiments are conducted on NVIDIA RTX A4000 GPUs.
Table 2. Scalibility analysis on Chengdu dataset.
Sampling Interval Rate 20% 40% 60% 80% 100%
[…]
$\mu=4$ minutes $\to\epsilon=15$ seconds MTrajRec 60.73 1048.0 63.58 968.5 64.73 929.4 65.39 914.6 65.79 904.4
T3s + Decoder 60.49 1098.9 63.32 968.7 64.23 961.9 65.04 942.3 65.60 926.3
T2vec + Decoder 62.17 966.1 64.92 946.1 64.95 972.3 65.31 946.8 66.51 915.2
RNTrajRec 60.80 998.0 62.85 931.3 64.19 909.1 65.46 835.0 67.66 886.0
MM-STGED 64.61 935.0 67.76 881.8 68.53 853.7 69.95 825.3 70.64 829.7
PTR 68.30 585.3 71.57 515.6 72.74 512.5 73.56 488.3 74.12 483.0
$\mu=2$ minutes $\to\epsilon=15$ seconds MTrajRec 69.85 919.8 72.58 907.4 73.97 902.3 74.32 891.3 74.52 885.7
T3s + Decoder 68.86 909.5 72.29 897.2 73.63 883.8 74.53 868.8 74.62 857.5
T2vec + Decoder 69.47 905.4 72.77 858.3 73.80 855.8 74.89 831.3 75.69 783.6
RNTrajRec 68.05 991.0 70.20 856.9 71.17 834.4 72.62 795.3 75.80 757.0
MM-STGED 71.65 865.0 75.32 773.0 75.49 752.1 76.94 747.7 78.14 696.0
PTR 76.29 452.8 79.37 411.5 80.37 395.4 81.18 377.4 81.76 366.2
$\mu=1$ minute $\to\epsilon=15$ seconds MTrajRec 75.39 937.5 78.53 835.1 80.01 794.3 80.93 725.7 81.12 718.4
T3s + Decoder 76.49 917.1 79.08 824.9 80.65 767.8 80.82 716.5 80.90 713.0
T2vec + Decoder 75.89 845.4 79.09 746.8 79.41 752.3 81.21 742.2 81.69 714.1
RNTrajRec 75.65 846.2 79.48 784.8 79.61 769.3 81.74 742.5 81.88 702.5
MM-STGED 76.02 857.7 79.86 734.7 82.25 676.4 83.44 663.0 84.26 633.5
PTR 81.62 397.5 84.59 352.0 85.75 328.7 86.73 305.4 87.17 290.8
5.5. Experimental Results (RQ1)
Table 1 reports a comparison of the results between our model and the baselines across various sampling intervals of sparse trajectories on both the Chengdu and Porto datasets. As the sampling interval increases, the map-matched trajectory recovery task is more difficult. Due to the more complex road network structure in Porto compared to Chengdu, with Porto having 100.9 roads / $km^{2}$ and Chengdu having 38.1 roads / $km^{2}$, the accuracy of road segment recovery is lesser in Porto than in Chengdu.
Compared to other baselines, our model achieves superior performance across all metrics, with an average improvement of 16.51% on the Chengdu dataset and 9.35% on the Porto dataset. Notably, it shows substantial gains in RMSE metrics, with an average reduction of 351.8 meters in Chengdu and 119.2 meters in Porto. This indicates PTR generalizes effectively, enabling it to accurately recover trajectories with varying sampling intervals. By fine-tuning at the specific sampling interval, PTR + FT, our model further improves Acc by 1.72% and 0.32%. These outstanding results can be attributed to the strong generalization capability of the PLM in processing sparse trajectories with varying sampling intervals. Additionally, the trajectory embedder plays a pivotal role in connecting trajectory data with the PLM, efficiently capturing spatial-temporal correlations and converting trajectory embeddings into a format that the PLM can effectively process and understand.
We also analyze the convergence rate of PTR across different sampling intervals, as shown in Figure 5 and  6. It demonstrates rapid convergence within just a few epochs on both the Chengdu and Porto datasets. While fine-tuning enhances performance further, the model stabilizes by the 3th epoch, indicating that PTR can deliver satisfactory results within a short training period.
5.6. Scalibility Analysis (RQ2)
To evaluate the effectiveness of PTR in trajectory recovery when dense trajectory data is limited, we conduct scalability experiments on the Chengdu dataset. Specifically, we train the model using subsets of the training set at 20%, 40%, 60%, 80%, and 100%, and evaluate performance on the test set. The results are presented in Table 2.
As the size of the training set increases, we observe performance improvements across all baseline models. Notably, with only 20% of the training data, PTR already surpasses most baselines, trailing only MM-STGED. With 40% of the training data, PTR outperforms all other models, demonstrating its strong generalization ability. This advantage underscores the practicality of PTR in real-world scenarios, where obtaining large-scale dense trajectory data is often challenging.
Table 3. Ablation study on two datasets with sampling intervals at 2 minutes.
[…]
Chengdu w/o explicit trajectory prompt 81.04 84.08 88.07 205.8 394.2
w/o implicit trajectory prompt 81.43 84.28 88.14 191.4 371.1
w/o trajectory prompt 80.70 83.61 87.73 225.9 407.4
w/o road network 81.28 84.22 88.04 194.5 375.6
w/o reference tokens 80.83 83.36 86.68 236.6 437.2
PTR 81.76 84.31 88.38 187.4 366.2
Porto w/o explicit trajectory prompt 66.13 66.19 81.98 157.9 314.6
w/o implicit trajectory prompt 66.18 66.22 82.01 149.6 298.1
w/o trajectory prompt 65.52 65.89 81.51 171.0 339.4
w/o road network 66.11 66.27 81.90 150.7 302.8
w/o reference tokens 65.89 65.07 80.57 165.4 327.7
PTR 66.40 66.52 82.14 141.9 294.6
5.7. Ablation Study (RQ3)
To evaluate the effects of various components in PTR, we conduct ablation experiments on the Chengdu and Porto datasets with sampling intervals of 2 minutes. We compare PTR with the following five variants.
  • w/o explicit trajectory prompt: We remove the explicit trajectory prompt.
  • w/o implicit trajectory prompt: We use the ‘[MASK]’ token in the BERT to represent the missing location.
  • w/o trajectory prompt: We remove both the explicit and implicit trajectory prompt, and the missing points are represented by the ‘[MASK]’ token in BERT.
  • w/o road network: We only use the Learnable Fourier Feature for encoding the observed trajectory points while removing the road network feature.
  • w/o reference tokens: We remove the trajectory feature transformation layer in the trajectory embedder module.
As shown in Table 3, each component of PTR is essential. The explicit trajectory prompt provides task and trajectory-related textual information, enabling PLMs to effectively understand the trajectory recovery tasks. Omitting this component leads to a performance decrease. When the implicit trajectory prompt is removed, many missing locations are replaced by the ‘[MASK]’ token in BERT, limiting the model’s ability to capture external factors for missing points, which results in suboptimal performance.
For the trajectory point embedder, excluding road network for observed points causes the model to rely solely on latitude and longitude, neglecting the relationship between trajectories and the road network, which demonstrates that road network is crucial for accurately modeling trajectory information. Additionally, removing reference tokens results in a performance decline, as PLMs are unable to directly interpret trajectory data.
6. Conclusion
In this paper, we propose PTR, a novel framework for trajectory recovery leveraging PLM. PTR efficiently handles trajectory recovery even with limited dense trajectory data, demonstrating strong scalability. The model is trained on datasets with various sparse trajectories, enabling it to effectively generalize to different sampling intervals. To model external factors, we introduce an implicit trajectory prompt to model road conditions, which provides rich information for missing trajectory points. We also develop a trajectory embedder to capture the spatial-temporal dependencies of sparse trajectories and convert them into a format interpretable by PLMs. Experimental results on two datasets with three different sampling intervals validate the efficacy and scalability of the proposed model.
Appendix A Dataset
Table 4. Dataset Description.
Types Chengdu Porto
Sampling interval 15s 15s
# Trajectory 118,354 322,079
# Road segment 2504 2224
Latitude range 30.655 $\sim$ 30.727 41.142 $\sim$ 41.174
Longitude range 104.043 $\sim$ 104.129 -8.652 $\sim$ -8.578
Length / Width 8.22 km / 8.00 km 6.19 km / 3.56 km
Appendix B Baseline Setting
We choose the following ten methods as baselines, including five free space trajectory recovery and five map-matched trajectory recovery.
B.1. Free-space Trajectory Recovery
Free space trajectory recovery first recovers trajectory points and then projects the trajectory onto the road network.
  • HMM (Newson and Krumm, 2009) + ShortestPath first projects the sparse trajectory on the road network based on the hidden markov model (HMM), and then calculate the shortest path.
  • Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009) linearly interpolates missing trajectory points and then implements HMM to perform the map matching process.
  • MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009) first divides the area of interest into grids to identify frequently traveled routes between sparse trajectory points. Then, it assumes that vehicle movement maintains a constant speed for trajectory recovery. Subsequently, it employs HMM to project the trajectory onto the road network.
  • DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009) incorporates a sequence-to-sequence framework and Kalman filtering to recover trajectory points. Subsequently, it applies an HMM to yield a trajectory that is constrained by the map.
  • AttnMove (Xia et al., 2021) + Rule uses attention to predict the missing road segments and uses the central location as the moving rate.
B.2. Map-matched Trajectory Recovery
Map-matched trajectory recovery can directly recover the trajectory on the road network.
  • MTrajRec (Ren et al., 2021) utilizes a sequence-to-sequence framework with Gated Recurrent Units (GRU) as the key component for trajectory recovery. It optimizes road segment and moving rate prediction through multi-task learning.
  • T2vec (Li et al., 2018) + Decoder is a deep learning model for trajectory similarity learning. We use its encoder to embed the sparse trajectory and use the decoder of the MTrajRec to recover the trajectory in an autoregressive manner.
  • T3s (Yang et al., 2021) + Decoder uses LSTM and attention mechanisms to encode sparse trajectory data effectively. Then, the MTrajRec decoder is utilized to recover the trajectory.
  • RNTrajRec (Chen et al., 2023b) leverages the Transformer to capture the spatial-temporal correlation of the sparse trajectories. It also takes into account the relation between the trajectory and the road network.
  • MM-STGED (Wei et al., 2024) models sparse trajectories from a graph perspective and recovers trajectories by capturing micro and macro semantic information.
Appendix C Evaluation Metrics
We adopt five widely used metrics to evaluate the effectiveness of our model, following previous works (Wei et al., 2024; Chen et al., 2023b; Ren et al., 2021). For road segment recovery, we use Accuracy (Acc), Recall, and Precision (Prec) to assess the alignment between the true road segments $\mathcal{E}_{p}=\{e_{1},\cdots,e_{m}\}$ and the predicted road segments ^$\hat{\mathcal{E}}_{p}=\{\hat{e}_{1},\cdots,\hat{e}_{m}\}$. A higher value in these metrics indicates a more accurate road segment recovery. The metrics are formally defined as follows:
(17) $\begin{split}\textbf{Acc}&=\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{e_{i}=\hat{e}_{i}\}\times 100\%,\\ \textbf{Recall}&=\frac{|\mathcal{E}_{p}\cap\hat{\mathcal{E}}_{p}|}{|\mathcal{E}_{p}|}\times 100\%,\\ \textbf{Prec}&=\frac{|\mathcal{E}_{p}\cap\hat{\mathcal{E}}_{p}|}{|\hat{\mathcal{E}}_{p}|}\times 100\%,\end{split}$
where $\mathbbm{1}\{\cdot\}$ is the indicator function, where $e_{i}=\hat{e}_{i},\mathbbm{1}\{e_{i}=\hat{e}_{i}\}=1$, elsewise $\mathbbm{1}\{e_{i}=\hat{e}_{i}\}=0$.
To evaluate the recovered GPS coordinates, we employ the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) to quantify the distance error between the true trajectory $\mathcal{T}_{m}={q_{1},\cdots,q_{|\mathcal{T}_{m}|}}$ and the predicted trajectory $\hat{\mathcal{T}}_{m}={\hat{q}_{1},\cdots,\hat{q}_{|\hat{\mathcal{T}}_{m}|}}$. The formulas for MAE and RMSE are given as follows:The formulas of MAE and RMSE are as follows:
(18) $\begin{split}\textbf{MAE}&=\frac{1}{|\mathcal{T}_{m}|}\sum_{i=1}^{|\mathcal{T}_{m}|}|\mathrm{RN\_dist}(q_{i},\hat{q}_{i})|,\\ \textbf{RMSE}&=\sqrt{\frac{1}{|\mathcal{T}_{m}|}\sum_{i=1}^{|\mathcal{T}_{m}|}|\mathrm{RN\_dist}(q_{i},\hat{q}_{i})|^{2}}\end{split}$
Here, following (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024), $\mathrm{RN\_dist}(q,\hat{q})$ signifies the shortest distance along the road network between the trajectory points $q$ and $\hat{q}$. Both MAE and RMSE are denoted in meters. Lower values of these metrics indicate a higher level of accuracy in the recovery results.
Appendix D Appendix Experiments
D.1. Covergence Rate
Figure 5. Convergence rate of PTR and fine-tuning for three sampling intervals on Chengdu.
Figure 6. Convergence rate of PTR and fine-tuning for three sampling intervals on Porto.
We evaluate the convergence speed of our model on the Porto and Chengdu datasets. As illustrated in Figure 5 and 6, the model reaches near-optimal performance by the 10th epoch. The trained model demonstrated robust accuracy across three sampling intervals (PTR: $u=1$ minute, PTR: $u=2$ minutes, and PTR: $u=4$ minutes), confirming that PTR can generalize effectively to sparse trajectories with varying sampling intervals. Furthermore, after fine-tuning at each sampling interval, the performance shows minimal change in performance by the 3rd epoch, further proving its fast convergence speed.
Table 5. Hyperparameter range and optimal value.
Parameter Range
The number of reference tokens $K$ 128, 256, 512, 1024
LoRA rank $r$ 4, 8, 16, 64
The loss function weight $\lambda$ 0.1, 1, 10, 100
Figure 7. Hyperparameter analysis of the number of reference tokens $K$ of trajectory feature transformation.
Figure 8. Hyperparameter analysis of the rank $r$ of LoRA.
Figure 9. Hyperparameter analysis of the weight $\lambda$ in the loss function.
D.2. Hyperparameter Study
To explore the impact of hyperparameters on model performance, we conduct hyperparameter analysis using the Chengdu and Porto datasets with the sparse trajectory sampling interval of 2 minutes. We consider three important hyperparameters, including the number of reference tokens $K$, the rank $r$ of LoRA, and the loss function weight $\lambda$. Details regarding the range of selection and the optimal values for these hyperparameters are presented in Table 5.
D.2.1. The Number of Reference Tokens $K$
We set the parameter $K$ from the set $\{128,256,512,1024\}$ to explore its impact on trajectory recovery. The experimental results are shown in Figure 7. As $K$ increases, the effectiveness of trajectory recovery also improves. This suggests that having a larger token space aids in accurately representing trajectory features. When $K$ is larger than 512, the performance improvement of PTR becomes marginal, while the computational cost will increase. To balance the performance and efficiency of the model, we set $K$ to 512.
D.2.2. The Rank $r$ of LoRA
As shown in Figure 8, we set the range of $r$ to 4, 8, 16, and 64 to explore its impact. It is observed that the model performs well when $r=4$. As $r$ increases, the accuracy will continue to increase, but it is not obvious. This suggests that PLMs encapsulate substantial domain expertise through training on extensive corpora, rendering them adaptable to trajectory recovery tasks with minor adjustments. Yet as $r$ increases, the number of parameters also increases, making the model training require more memory. Therefore, we set $r=8$ to balance the performance and resources.
D.2.3. The Loss Weight $\lambda$
As shown in Figure 9, as the loss function weight $\lambda$ increases, the model performance first improves and then decreases. This is because a smaller $\lambda$ will make the model focus on road segment recovery, while a larger $\lambda$ will focus on moving rate recovery. To balance these two tasks, we set the value of $\lambda$ to 10.
Figure 10. Case study on the Chengdu dataset. Red points represent the truth trajectory points and blue points represent the recovered trajectory points.
D.3. Case Study
We conduct a case study on the Chengdu dataset to visualize the trajectory recovery performance with various baselines. As shown in Figure 10, we draw the truth and recovered trajectory, where red points represent the truth trajectory points and blue points indicate the recovered trajectory points. We find that the recovered trajectory aligns well with the road network, demonstrating the effectiveness of using road segment and moving rate to represent trajectory point. To facilitate a more intuitive comparison of recovery performance among different models, we focus on two distinct regions, labeled as A and B, for visual analysis. In region A, characterized by a relatively simple road network structure, all models can accurately recover road segments. Among them, PTR maps trajectory points better by leveraging PLMs. In contrast, region B exhibits a more intricate road network with multiple accessible routes. RNTrajRec captures spatial-temporal correlations of trajectories that are not adequate and recovers incorrect road segments. PTR not only excels in road segment recovery but also in accurately matching the actual trajectory points. This success can be attributed to its ability to model missing trajectory points through the implicit trajectory prompt, thereby introducing more valuable information and improving performance.
D.4. Computational Cost
We use an NVIDIA RTX A4000 card to conduct cost analysis on the Chengdu and Porto datasets with a 2-minute sampling interval by considering model size, training time, and inference time. As shown in Table 6, compared to the end-to-end deep learning model, PTR has the largest number of parameters due to its utilization of a pre-trained language model based on the BERT-small framework. Despite having a greater number of parameters, PTR exhibits quicker training and testing times as it can output results for all trajectory points at once, and the majority of parameters are frozen, unlike other models that necessitate autoregressive generation. It significantly accelerates both model training and inference processes.
Table 6. Computation Cost.
Dataset Chengdu / Porto
Methods Model size Train time Inference time
(MBytes) (min/epoch) (min)
DHTR 19 / 18 11.30 / 42.55 1.13 / 2.35
MTrajRec 49 / 44 37.19 / 142.12 12.57 / 28.20
RNTrajRec 112 / 109 72.30 / 153.29 29.14 / 43.58
MM-STGED 34 / 31 41.27 / 138.33 16.27 / 30.58
PTR 176 / 175 24.18 / 54.14 2.56 / 10.38